OpenStudy (anonymous):

Solve the equation y + 2 = 6 for y. Please show your work and solution in the space provided. Solve the equation y + 2 = 6 for y. Please show your work and solution in the space provided. @Mathematics

OpenStudy (diyadiya):

y=6-2 y=4

OpenStudy (anonymous):

y=6-2 y=4